OpenStudy (anonymous):

Given: ABC is similar to XYC. Enter answer as a simplified, improper fraction. If BY = 4, YC = 7, XC = 10, then AC =

OpenStudy (anonymous):

u have picture?

OpenStudy (anonymous):

yeah but it does not come out on the screen its a flash picture do you think you could help me

OpenStudy (anonymous):

look on the botom click attach file

OpenStudy (anonymous):

i cant its not a picture its a question

OpenStudy (anonymous):

I be back

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

\[BC \approx YC\]\[AB \approx YX\]\[CX \approx AC \]\[\frac{7}{AC}=\frac{7}{10}\]

OpenStudy (anonymous):

crosss and time 7*10=7 70/7=10

OpenStudy (anonymous):

its not right :/

OpenStudy (anonymous):

yes, if u have picture I will do better